If a represents Ang's age five years from now, which of the following represents his age 17 years ago?
A. a-22
B. a-12
C. a+12
D. a-17
E. a+22

Found 2 solutions by sheldonbbtrocks, jim_thompson5910:
Answer by sheldonbbtrocks(115)

5 years from now-17 years ago=-12 years

D. a-12

Answer by jim_thompson5910(35256)

a = age five years from now (ie his future age in 5 years)

So his age right now is a-5 years old

His age 17 years ago is (a-5)-17 = a-5-17 = a-22 years old which is choice A